OpenStudy (anonymous):

Snookers Lumber can convert logs into either lumber or plywood. In a given day, the mill turns out three times Snookers Lumber can convert logs into either lumber or plywood. In a given day, the mill turns out three times as many units of plywood s lumber. It makes a profit of $25 on a unit of lumber and $40 on a unit of plywood. How many of each unit must be produced and sold in order to make a profit of $15080?

OpenStudy (anonymous):

3*40 +25= 145 15080/145 = 104 lumber = 104 units ply = 104*3 = 312 units

OpenStudy (anonymous):

what is the star?

OpenStudy (anonymous):

multiplication